The owners equity of Goodman Interiors appears on the balance sheet as 300,000 and is equal to one-third of the total assets. Compute the amount of total liabilities.
Answer #1
Let Total Asset= X
Owner Equity = 1/3 of Total Asset
Owner Equity = 1/3*X
300000=1/3*X
Total Asset =900000
Total Liabilities= Total Asset- Equity
=900000-300000= $600000